Project Narrative

Experimental Medium Energy Physics

at Arizona State University

September 2021

Project Title:  Experimental Medium Energy Physics at Arizona State University
Organization:  Arizona State University, College of Integrated Sciences and Arts

7271 E. Sonoran Arroyo Mall, Mesa, AZ 85212-6415

Principal Investigator: Michael R. Dugger [(480) 965-0728, dugger@jlab.org]
ASU Admin. Contact: Katrina Smith [(480) 965-4225, Katrina.Smith@asu.edu]
Submitted to:  Department of Energy, Office of Nuclear Physics
Funding Opportunity: DE-FOA-0001820
DOE Contact:  Gulshan Rai
Research Area:  Medium Energy
New Grant Period:  9/1/2022 - 9/1/2025

Contents

1 Introduction 1

2 Research to be undertaken under proposed award 3 2.1 The CLAS12 detector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.2 The GlueX detector in Hall D . . . . . . . . . . . . . . . . . . . . . . . . . . 4

2.2.1 TPOL: the ASU triplet polarimeter . . . . . . . . . . . . . . . . . . . 5 2.3 Probing the Ξ baryon spectra . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.4 Exotic hadrons . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

2.4.1 ASU analysis of GlueX K^{∗}K data . . . . . . . . . . . . . . . . . . . . 8

3 Past accomplishments and works in progress 8

4 Research plan 8

5 Impact of other work on this grant’s resources 9

6 Roles and responsibilities 9

7 Human resource development 10

8 Requested resources and justification 10 8.1 Senior personnel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 8.2 Other Personnel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

8.2.1 Graduate Research Associates . . . . . . . . . . . . . . . . . . . . . . 10 8.2.2 Undergraduate students . . . . . . . . . . . . . . . . . . . . . . . . . 11

8.3 Fringe benefits. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 8.4 Total salaries, wages and fringe benefits . . . . . . . . . . . . . . . . . . . . . 11 8.5 Equipment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 8.6 Travel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 8.7 Other direct costs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 8.8 Total direct costs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 8.9 Indirect costs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 8.10 Total Direct and Indirect costs . . . . . . . . . . . . . . . . . . . . . . . . . . 12

9 Institutional support 12

10 Summary of meaningful metrics 13

Appendices 14

A Data management plan 14

B Bibliography and references cited 15 C Principal collaborators 16

C.1 Close collaborators . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 C.2 Advisors and postdoctoral sponsors of the investigators . . . . . . . . . . . . 16 C.3 Former post-doctoral participants in academic and lab positions . . . . . . . 16 C.4 Former Ph.D. students in academic and lab positions . . . . . . . . . . . . . 16

D Biographical sketches 17 D.1 Michael R. Dugger . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

Abstract

Using the experimental facilities of Jefferson Lab in Halls B and D, the Arizona State University (ASU) Meson Physics Group will conduct experiments to search for and characterize (1) excited baryons containing two or more strange quarks and (2) hybrid meson states. Both these types of particles have been specifically predicted by hadron models based on quantum chromodynamics (QCD), exploiting flavor symmetry and gluonic excitations. The detailed knowledge of the spectrum of these states that will be obtained is essential for meeting the goals of NSAC Long-Range Plan.

1 Introduction

The ASU Meson Physics Group is primarily interested in determining the existance and properties of hadronic resonance states. We have chosen to focus our efforts on resonance states that involve strangeness. In particular, we have concentrated on the excited Ξ baryons and the meson resonances that decay KKπ.

As stated in the current NSAC Long-Range Plan [1]: “[t]he new capabilities of the 12-GeV era facilitate a detailed study of baryons containing two and three strange quarks. Knowledge of the spectrum of these states will further enhance our understanding of the manifestation of QCD in the three-quark arena.” It is our goal to provide the detailed data regarding the excited Ξ baryons.

Improving our understanding of quark confinement in hadrons requires deepening our insight into the gluonic fields within those particles. Exploration of the spectrum and properties of exotic hybrid mesons has emerged as a key path forward for exploring quark confinement. Indeed, that topic was a major motivation for the upgrade of JLab, and is specifically mentioned multiple times in the NSAC Long-Range Plan [1].

For our study of mesonic resonances, the primary focus is to assist in fully mapping out the light meson nonets while pushing upwards in mass towards potential discoveries of exotic hybrid meson states that decay K^{∗}K. The presence of non-standard J^{P C} quantum numbers would be sufficient to clearly identify a meson as being “exotic”. To assist in the J^{P C} determination, the degree of photon polarization is very useful. The knowledge of the degree of photon polarization is required for determining beam asymmetries, which in turn, help establish the J^{P C} quantum numbers.

The ASU Meson Physics Group successfully proposed, designed, built, and now operates a first-of-its-kind triplet-photoproduction photon-polarimeter (TPOL) to measure the degree of linear polarization of the Hall D photon beam. The group will remain responsible for determining the polarization of the photon beam for all experiments in Hall D.

During the previous award period, undergraduate research was utilized to good effect. The undergraduate participants Patrick Walker, Rebecca Osar and Emily Lamagna worked on CLAS12 data. Mr. Walker analyzed K^{∗} K final states, successfully defended his honors thesis, graduated with his B.S degree and won the ASU Department of Physics Research Award. Ms. Osar has been analyzing the photoproduction of K Λ states from the proton. Ms. Lamagna is in the process of analyzing the reaction ep → epϕπ. Additionally, undergraduate student Anna Costelle won the ASU Women in Physics Award for Undergraduate Research for her presentation describing the development of an event generator for strangeness states. In May of 2020, undergraduate student Mohamed Mohamed won the College of Integrative Sciences and Arts (CISA) Undergraduate Reserach Award for his work on energy deposition studies using the Monte Carlo simulation software GEANT4 and for a study on the use of decision trees to assist in particle identification. Undergraduate student Kevin Scheuer presented his study of machine learning methods for Kπ at the spring 2020 CISA research symposium [?].

While we are very proud of our undergraduates, the primary focus of our work is accomplished by our graduate students. Graduate student Sebastian Cole (now Dr. Cole) successfully defended his dissertation that was based on studies utilizing AmpTools (partial wave analysis software) to find the J = 0, J = 1 and J = 2 contributions to the mass spectrum of K^{∗} K states in the mass region below 1.6 GeV. Sebastian won the CISA outstanding graduate student award in December of 2020 and I have nominated Dr. Cole for the Four Corners Section of the American Physical Society, Harry Lustig Award. Graduate student Brandon Sumner has refined his preliminary cross sections for γp → K^{+}K^{+}Ξ^{−}(1530), where Ξ^{−}(1530) → Ξ^{−}π^{0}. Brandon has also begun analyzing data for excited cascades to decay Ξ^{0}π^{−} as well as to K^{−}Λ. Graduate student Alan Gardner is starting his second year and assisted Sebastian Cole in building code used in the partial wave analysis of the K^{∗} K from GlueX data. Prior to his graduate studies, Alan conducted a preliminary search for the Ξ^{−} baryon using CLAS12 data.

The PI performed studies of systematics of the polarization measurement from the ASU triplet polarimeter associated with the converter thickness. Performed final-stage analysis of polarimeter data for the 2020 data set (ongoing). Assisted students in analysis projects related to GlueX and CLAS12 data.

This proposal seeks support for a program of measurements that will provide improvements in our knowledge of hadron spectroscopy, focused on activities at Jefferson Lab (JLab)

(1) within Hall B using the CEBAF Large Acceptance Spectrometer (CLAS) newly upgraded for experiments using the 12 GeV beam at JLab (CLAS12) and (2) within Hall D using the GlueX detector.

We will obtain and analyze data using the GlueX and CLAS12 detectors to explore the spectrum of multi-strange-quark baryons in order to assist in resolving the missing resonance problem for Ξ states and to refine our knowledge of the excited Ξ properties. We will also map out the light mesons that decay KKπ with a particular interest in finding meson resonances that exhibit exotic J^{P C} quantum numbers.

Currently, the ASU Meson Physics Group is comprised of the principal investigator Associate Professor Michael Dugger, Graduate Research Assistant Brandon Sumner, Graduate Research Assistant Alan Gardner, senior undergraduate student Anna Costelle, senior undergraduate student Emily Lamagna and junior undergraduate student Rebecca Osar. The proposal seeks steady-state funding for supporting the principal investigator (PI) and two graduate students.

2 Research to be undertaken under proposed award

Prior to any analysis performed to uncover hadronic states within physical data, that data must be aquired using specially designed, built and operated detector systems. Thus, it is natural to begin the discussion of the proposed research by providing brief overviews of those detector systems used in measuring the data that we further analyze.

2.1 The CLAS12 detector

While the CLAS12 detector is designed primarily to measure deeply exclusive scattering using electron beams, a forward tagger has been installed within CLAS12 that enables quasi-real photon experiments with q^{2} as low as 0.01 GeV^{2}. The ASU efforts in Hall B make use of this forward tagger to perform photoproduction measurements from the cryogenic hydrogen target within the CLAS12 detector.

As seen in Fig. 1, CLAS12 is composed of two main detector arrays: the Central Detector (CD) and the Forward Detector (FD). The FD covers the range of 5^{∘} to 40^{∘} in polar angle, whereas the CD covers the polar angles from 40^{∘} to 135^{∘}. The FD is expected to provide separation of identified K^{±} and π^{±} having momenta less than 3 GeV/c, and separation of protons and K^{±} with momenta less than 4.5 GeV/c.

Figure 1: Schematic diagram for CLAS12. The labeled subsystems are described in the text. A cryogenic target cell is located at the center of the cylindrically-symmetric CD. Surrounding the cryogenic target cell is a Silicon Vertex Tracker (SVT) yielding 1 mrad angular resolution in azimuthal angle and 15 mrad in polar angle for charged tracks, as well as providing vertex resolution better than 2 mm. Outside the SVT is the cylindricallysymmetric Central Time-of-Flight system (CTOF). The CTOF scintillator array has a timing resolution of 50 ps. The Central Electromagnetic Calorimeter forms the next cylindrical layer, consisting of a scintillating-fiber/tungsten-powder array. These detectors are housed within a 5-T superconducting solenoid magnet of aperture 0.78 m. The original CLAS superconducting toroidal magnet has been repurposed for CLAS12, with the same six-sector azimuthal symmetry as the original detector.

Downstream of the SVT are the High Threshold Cherenkov Counters and the Forward Tagger (FT) arrays. Electrons that scatter with angles between 2.5^{∘} and 4.5^{∘} are detected in the FT, enabling quasi-real photon experiments such as those we will undertake in this work. Beyond the FT are three regions of drift chambers, forming the Forward Tracking Chamber (FTC). The chambers are arranged in a fashion similar to the original CLAS design, with Region 1 drift chambers located before the torus magnet, Region 2 within the torus magnet, and Region 3 behind the torus. Ultimately, the spatial resolution for this chamber package is expected to be a factor of 3 better than the original CLAS system. Downstream of the FTC are the Low Threshold Cherenkov Counters (LTCC), which are modified versions of the original CLAS Cherenkov counters with the same external geometry. As with the LTCC, the Forward Time-of-Flight (FTOF) system has the same geometry as the original CLAS system, but has been upgraded to obtain 80-ps timing resolution. The detector furthest downstream is the Forward Electromagnetic Calorimeter, an array retained and repurposed from the original CLAS detector.

2.2 The GlueX detector in Hall D

The Hall D GlueX detector was built to search for gluonic excitations within mesons with masses up to 2.5 GeV [10, 11]. Distinguishing states with gluonic excitations from other meson states will require the determination of exotic quantum numbers, J^{P C} that are unlike the assignments possible with traditional constituent quark models of meson states. A full partial wave analysis will be needed to obtain J^{P C} for the observed states. To minimize possible ambiguities in obtaining the required J^{P C} information from the partial wave analysis, all decay products from the candidate hybrid meson must be identified with good resolution. The identification of all decay products is greatly enhanced when the detector has a full acceptance in decay angles. For these reasons the GlueX detector (Fig. 2) was designed to have 4π solid angle coverage, with nearly full hermeticity, and good momentum and energy resolution of photons and charged particles.

Figure 2: Diagram of Hall D and the GlueX detector. Electrons with energies of 12 GeV incident on a thin (50 or 20 μm) diamond radiator produce a beam of linearly polarized photons (≈40% polarization) with E_{γ} = 8.5-9 GeV. The tagging of the postbremsstrahlung electron determine the photon energy to within 0.1%. The photons strike a 30-cm-long liquid hydrogen target surrounded by a start counter. Past the start counter, the next detector elements are cylindrical driftchambers, which are surrounded by a cylindrical electromagnetic calorimeter made of lead and scintillating fibers. The target, start counter, drift chambers, and electromagnetic calorimeters are all contained within a refurbished 2.25 T superconducting solenoid that provides the momentum-analyzing magnetic field. On the downstream side of the solenoid is a time-of-flight system and a 3000-element lead-glass electromagnetic calorimeter. For 9 GeV incident photons, the GlueX detector has 4π acceptance for photons and charged particles to facilitate the partial-wave analyses of many-particle final states. The acceptance of final-state particles is typically better than 95%, and is very uniform over the detector. Moreover, the near hermetic design of the detector is well suited for determining the masses and J^{P C} for mesons with masses from 1.5 to 2.5 GeV/c^{2}.

For increased precision in kaon particle identification, the newly acquired DIRC (Detection of Internally Reflected Cherenkov) is placed just in front of the FCAL. The DIRC has as the main component four boxes filled with fused silica bars from the BaBar experiment. When installed in the BaBar experiment, these boxes provided good kaon separation up to 4 GeV [12]. Similar performance is expected when utilized for the GlueX experiment, greatly improving kaon identification, which will facilitate many analyses including those ASU is particularly targeting.

2.2.1 TPOL: the ASU triplet polarimeter

The partial-wave analyses to be undertaken in the GlueX search for hybrid mesons require an absolute uncertainty of ±0.04 in the photon beam polarization. In order to obtain that goal, the ASU Meson Physics Group designed and constructed a detector system (called TPOL) that measures photon beam asymmetries with this triplet production process (pair production off a recoiling atomic electron), and installed the system in the Hall D beamline.

Figure 3: Interior of TPOL vacuum chamber with SSD and production target ladder assembly (from beam left).

Figure 4: TPOL vacuum chamber and electronics housings installed in Hall D beamline (from beam right).

TPOL uses a double-sided silicon strip detector (SSD) manufactured by Micron Semiconductor to measure timing, energy deposition, and the azimuthal distribution of recoil electrons coming from the triplet production process. Fig. 3 shows the interior of the TPOL vacuum chamber; the beam transits that figure from right to left. The SSD has 32 azimuthal sectors on the ohmic side and 24 concentric rings on the junction side, with an outer active diameter of 70 mm and an inner active diameter of 22 mm.

The TPOL installation in the Hall D beamline is shown in Fig. 4, seen from beam right. The housings consist of three boxes attached to one another. The largest rearmost box in the figure is the vacuum chamber, the middle box is the preamp enclosure, and the smallest box is the distribution box that routes voltages and signals to and from the preamps. Cables attached to the distribution box include the signal lines (white), high voltage bias (red) and low voltage (gray) lines.

To determine the degree of photon beam polarization P in practice, the value of B obtained from fitting the azimuthal distribution to A [1 + B cos(2ϕ)] is divided by the analyzing power Σ.

ASU is positioned to continue providing the important polarization information for all Hall D experiments with TPOL during the proposed award period.

2.3 Probing the Ξ baryon spectra

Studies of two and three-strange-quark baryons have great potential to enhance our understanding of the QCD structure within baryons. The “computational cost” for predictions using lattice QCD are greatly influenced by the number of light quarks (u and d) present in the baryon studied. For that reason, studies that extend our knowledge of the properties of states with two strange quarks should be extremely useful in deepening insight into how all baryons are assembled from quarks.

Table 1: “Well-known” Ξ resonances listed in the PDG summary [2]. State J^{P} Width PDG (MeV) rating Ξ(1318) \frac{1}{2}^{+} **** Ξ(1530) \frac{3}{2}^{+} 9.1±0.5 **** Ξ(1690) \frac{1}{2}^{−}? 10±6 *** Ξ(1820) \frac{3}{2}^{−} 24±6 *** Ξ(1950) ? 60±20 *** Ξ(2030) ≥ \frac{5}{2}? 21±6 *** The ASU group is participating in experiments which will provide the largest-ever sample of photoproduced cascade Ξ baryons. Using the CLAS12 and GlueX detectors, our measurements should greatly enhance knowledge of the excitations of these baryons, which in turn will provide an empirical “proving ground” for QCD-based descriptions of baryons with varying numbers of heavy quarks.

While the “missing resonance” problem for the nucleon system is much discussed, the missing resonance problem is even more striking when looking at the Ξ sector. For example, the Isgur and Capstick relativistic constituent quark model [3] predicts 44 Ξ states with masses below 2.5 GeV, yet the PDG [2] only lists six cascade resonances with ratings of three or four stars, as indicated in Table 1.

Table 2: PDG values for branching ratios of excited Ξ states [2]. State ΛK ΣK Ξπ Ξ(1530) 100 % Ξ(1690) seen seen seen Ξ(1820) large small small Ξ(1950) seen seen? seen Ξ(2030) 20% 80% small Looking more closely at the Ξ states, photoproduction cross sections for Ξ resonance states are observed to be much smaller than for nucleon resonances. However, the widths of the Ξ excitations are much narrower than those for the nucleon, which should make them easier to identify [4]. As seen in Table 1, the well-known cascade resonances also have mass separations that are greater than the resonance widths, which again should make identification of excited Ξ states far easier than for the N^{∗} and Δ states (with widths of ∼150 MeV). Even with those advantages, however, no new cascade states have been discovered in the last 20 years. The 1989 discovery of Ξ(1690) [5] was plagued by backgrounds from a Σ^{−}(C, Cu) → Ξ^{−}π^{+} inclusive reaction. The large solid angle of the CLAS12 detector (discussed further in the next subsection) will permit using exclusive reactions to markedly reduce backgrounds. Moreover, the CLAS12 detector will simultaneously measure Ξ^{∗} decay modes, allowing for the determination of branching ratios that are, as shown in Table 2, not well determined for most of the three and four star Ξ^{∗} states [2]. Of particular interest is the apparent suppression of the excited cascades going to Ξπ. For the Ξ(1530), the branch to Ξπ is given in the PDG as 100%, but for all other cascades, the observed decay branches are merely characterized as “small” or “seen”. Measuring the Ξ^{∗} branching ratios for states that decouple from Ξπ can help shed light on why the Ξ^{∗} have such narrow widths.

The information obtained on mass splittings within the Ξ system will be extremely useful in testing lattice QCD and QCD-inspired models of these baryons. At present, the only baryonic mass splittings available come from the ground state octet and decuplets (e.g. n-p splitting or ground state Δ splitting). All the data will help address the issue of how quarks and gluons assemble into the hadrons observed in nature.

2.4 Exotic hadrons

To deepen understanding of how confinement arises, ASU is working with the GlueX Collaboration in Hall D to search for exotic hybrids with masses up to 2.5 GeV. Exotic hadrons (“exotics”) are classified as one of three types: glueballs, tetraquarks, and hybrids. Glueballs are states that completely lack valence quarks. The lightest glueballs should have J^{P C} = 0^{++} or 2^{++}, and are expected to mix with constituent quark states with the same quantum numbers. Tetraquarks are comprised of two q\overline{q} states, and are the product of two SU(3) octets that can be represented as

8 ⊗ 8 = 27 ⊕ 10 ⊕ \overline{10} ⊕ 8_{1} ⊕ 8_{2} ⊕ 1.

Hybrid states in a constituent q\overline{q}g model are built from the usual q\overline{q} constituent-quark representation with the promotion of one or more gluons from virtual to real. For a hybrid state, a q\overline{q} colored state interacts with real constituent colored gluon(s) to form an overall colorless state. In contrast to glueballs and tetraquarks, hybrids are only expected to exist in nonets similar to those for the constituent quark model. Exotics, however, can have quantum numbers J^{P C} that differ from those of hadrons constructed with the simple constituent quark model. Finding particles with exotic quantum numbers J^{P C} serves as the “smoking gun” for determining if a state is exotic.

Some guidance in the search for hybrid states has come from lattice QCD predictions. For example, a lattice calculation performed by the Hadron Spectrum Collaboration (HSC) in 2013 [6] found hybrid states in supermultiplets with J^{P C} = 0^{+−}, 1^{−+}, 2^{+−}, and 1^{−−}. This HSC model description suggests that those hybrids have the structure ([q\overline{q}]_{8_{c}} ⊗ G_{8_{c}}). States of particular importance for the search of hybrids have the exotic quantum numbers of J^{P C} = 0^{+−}, 1^{−+}, 2^{+−}.

Table 3 shows several possible decays of exotic mesons of interest for the GlueX experiment. Currently, there are two exotic J^{P C} = 1^{−+} states (π_{1}(1400) and π_{1}(1600)) listed by the Particle Data Group [2]. For the π_{1}(1400), the decay modes that have been seen are ηπ^{0} and ηπ^{−}, while for π_{1}(1600) the decays seen are b_{1}π, η′π^{−} and f_{1}π. The interpretation of the π_{1}(1400) as a hybrid is problematic [7–9]. A hybrid meson (q\overline{q}g) state can either have an angular excitation between the quark+anti-quark system and the gluon (said to be “gluon excited”) or the excitation can be exclusively between the quark and anti-quark (“quark excited”). For the gluonic angular-excitation, within the q\overline{q}g constituent model, a decay of a hybrid into two fundamental mesons is suppressed. This would make the π(1400) → ηπ decay unlikely in the gluon-excited case. However, a quark excitation can proceed into two fundamental mesons, though the branch into the η-singlet is favored over the η-octet. Since no search for π(1400) → η′π has been successful, there is some doubt that the π_{1}(1400) is a hybrid. Instead, some have suggested that the π_{1}(1400) is a tetraquark state [7–9].

Table 3: Possible decay modes of exotic hybrids that will be explored by GlueX.

Hybrid  Reachable with  Reachable with  Challenging even
meson J^{P C} limited running  longer running  with longer running
b_{0}  0^{+−}  —  —  π(1300)π, h_{1}π,
f_{1}π, b_{1}π
h_{0}  0^{+−}  —  b_{1}π  h_{1}η
h′_{0}  0^{+−}  —  K_{1}(1270)K  K(1460)K, h_{1}η
π_{1}  1^{−+}  πρ  πb_{1}, πf_{1}, πη′  ηa_{1}
η_{1}  1^{−+}  ηf_{2}, a_{2}π  ηf_{1}, ηη′  π(1300)π, a_{1}π
η′_{1}  1^{−+}  K^{∗}K  K_{1}(1270)K, K_{1}(1410)K,  —
K_{2}^{∗}K, ηη′
b_{2}  2^{+−} ωπ, a_{2}π, ρη  —  f_{1}ρ, a_{1}π, h_{1}π, b_{1}η
h_{2}  2^{+−}  ρπ  b_{1}π, ωη  f_{1}ω
h′_{2}  2^{+−}  —  K_{1}(1270)K, K_{1}(1410)K  f_{1}ϕ
K_{2}^{∗}K, ϕη

In addition to the π_{1} exotic states, there should be corresponding isospin I = 0 nonet members if the π_{1} is a hybrid. Establishing these isoscalar members for the nonets is important in order to verify these exotics as being hybrids. The tetraquark state with J^{P C} = 1^{−+} excludes the presence of an isosinglet, and is fully contained within the 20-plet (10 ⊕ \overline{10}). Thus, identifying an isosinglet state with J^{P C} = 1^{−+} excludes that state as being a tetraquark.

2.4.1 ASU analysis of GlueX K^{∗}K data

To illustrate the types of analysis we can currently perform with GlueX data, we discuss here portions of the analysis of multi-kaon states undertaken by ASU graduate student Sebastian Cole

3 Past accomplishments and works in progress

4 Research plan

The timeline given here represents a “good faith” estimate that assumes that the accelerator schedule at JLab will proceed as planned, the requested funds for the proposed work are awarded at the level requested and no unforseen issues arise. The entire effort of the ASU Meson Physics Group is dedicated to this proposal and the work will proceed as rapidly as reasonably possible.

Throughout grant period, additional data will be taken with both Halls B and D and we will contribute towards the normal collaborative service work expected from us. For the GlueX collaboration in Hall D, we will continue our responsibility for maintaining the ASU-built triplet polarimeter. We will also continue analyzing TPOL data required for providing the GlueX collaboration the measured values of polarization for the photon beam that is incident on the GlueX target. Opportunistic physics analyses likely will present themselves as these analyses proceed, which may spawn projects for undergraduate and graduates students. As analyses progress, preliminary data will be presented at workshops and meetings. We will be deeply engaged in multiple analyses from both Hall B and GlueX running, will have mature software tools for analyzing data from those facilities, and be able to fold additional students into the ASU efforts.

During Year 1, graduate student Alan Gardner will assume on-site responsibilities for the triplet polarimeter located in Hall D of Jefferson Lab. Along with the hardware responsibilities, Alan will push forward with his partial wave analysis of resonant mesons with K^{+} K^{−} π^{0} decay products. A second graduate student funded by the new award will enter the group to extend the cascade analysis from, then, former student and newly minted PhD Brandon Sumner, to include data from both Halls B and D. In addition to the cascade work, the second graduate student will also assist with our CLAS12 analysis projects being conducted by our undergraduate students.

During Year 2, Alan Gardner will return to ASU and the second graduate student will assuming the on-site responsibilities for the triplet polarimeter.

During Year 3, The second graduate student will return from Jefferson Lab.

5 Impact of other work on this grant’s resources

PI Dugger is a tenure-track faculty member at ASU with an academic year contract. ASU requires all tenured and tenure-track faculty members to contribute effort during the academic year in the areas of teaching, research, and service. The PI will fully commit his entire research effort during the academic year to the research to be undertaken in this proposal, and all effort during the summer months will be allocated to this proposal. The PI has no effort committed to any other proposal.

6 Roles and responsibilities

The ASU Meson Physics Group is a relatively small research group with the PI involved and cognizant of work undertaken in all analyses related to the work discussed in this proposal. PI Dugger will be responsible for leadership of the analyses and publications related to the Hall B/CLAS12 and Hall D/GlueX data. He will also serve as the mentor for graduate students working on GlueX analyses, as well as undergraduate students working on the CLAS12 analyses. The PI will be responsible for maintaining the TPOL, manning shifts, and in the mentoring of students. Weekly group meetings will assist in coordinating and synchronizing all efforts and to provide “cross-training” for the students working with the group.

7 Human resource development

Based at one of the nation’s largest and most diverse major public research universities, instruction and mentoring form a central feature of this project. Mentoring and training of students is an essential feature of this proposal.

Among the undergraduate students who have worked with the PI, Robert J. Lee, was mentored by the PI, and graduated in Spring 2018 and has begun graduate studies in physics at the University of California at San Diego. Robert was an exceptional student that the PI worked closely with for three years. His honors thesis was on the analysis of FROST (Frozen Spin target) data from JLab CLAS data. He analyzed the spin observables associated linearly polarized photons incident on transversely polarized protons for the reaction γp → π^{+}n. Currently, The PI has a sophomore undergraduate student,

As expected, the scientific results of the investigations outlined in Section 2 will be broadly disseminated through national and international conference presentations, workshops, reports, and refereed publications, and will be archived in public databases.

8 Requested resources and justification

The justifications for the items listed in the proposal budget are provided here.

8.1 Senior personnel

Summer salary support is being sought for the Principal Investigator to enable his effort to continue during the summer months when his full time can be devoted to the research described above. The PI has extensive experience working in both Halls B and D at Jefferson Lab. While active in all areas of the project mentoring the graduate student and undergraduate students working on the hadronic production efforts in Halls B and D.

8.2 Other Personnel

8.2.1 Graduate Research Associates

Support is sought for a steady-state group size with two graduate students. The funding requested encompasses for each supported student 4.5 academic months per academic year and 1.5 summer months per calendar year. At ASU, this translates to 20 hours/week during the academic year and summer, which is considered 100% effort for a graduate student. As a graduate student completes his program at ASU, a new graduate student will assume the vacated line. Student participation will greatly enhance the productivity of the group beyond the efforts of the investigators, and also deliver the human resource benefits noted above, enhancing the nation’s supply of well-trained scientists. Graduate student salaries listed in the budget are based on the standard rate for graduate research associates in the Department of Physics with an additional three months of summer support.

8.2.2 Undergraduate students

Some travel????

8.3 Fringe benefits.

Arizona State University defines fringe benefits as direct costs, estimates benefits as a standard percent of salary applied uniformly to all types of sponsored activities, and charges benefits to sponsors in accordance with the federally-negotiated rates in effect at the time salaries are incurred. Benefit costs are expected to increase approximately 3% per year; the rates used in the proposal budget are based on the current Federally-Negotiated Rate Agreement plus annual escalation for out-years. The proposal’s Fringe Benefits were negotiated and approved by the U.S. Department of Health and Human Services. For purposes of this proposal the following rates are being used:

• Faculty members: 28.12%, 28.96%, and 29.83% of salary for Year 1, 2, and 3, respectively.

• Graduate research assistants/associates: 7.31%, 7.53%, and 7.76% of salaries for Year 1, 2, and 3, respectively.

• Undergraduate students (hourly): 2.16%, 2.22%, and 2.29% of wages for Year 1, 2, and 3, respectively.

8.4 Total salaries, wages and fringe benefits

A cost-of-living increase of 3% has been calculated in years 2 – 3 for each year of this project. The total request for salaries, wages and fringe benefits: Year 1 -$105,981, Year 2 -$109,632, Year 3 -$113,423.

8.5 Equipment

No funds are requested for equipment.

8.6 Travel

The travel budget was estimated in accordance with the University’s travel policy based on current air fares, current federal and ASU authorized per diem rates, airport shuttle services, and, if applicable, conference registration fees and/or car rental. At the time of this proposal, ASU’s travel system software provider, Concur Technologies, assesses a minimal charge per person for each travel expense report submitted. This expense is a direct cost charged per trip.

For each year of support, we have allocated six trips to Jefferson Lab, with four of those trips having a duration of ten days and nine nights, and the remaining two JLab trips having durations of six days and 5 nights. These trips will enable at least one additional person from ASU beyond the on-site graduate student(s) to participate in collaboration activities, meet service obligations, discuss analyses and results with collaborators, and make presentations at collaboration meetings. In addition to the JLab trips, we have allocated two trips per year for giving talks at conferences. The travel funds being requested: Year 1 -$15,902, Year 2 -$16,379, Year 3 -$16,870.

8.7 Other direct costs

Funds for \underline{materials and supplies} are requested at the level of approximately $4,000 per year for computing equipment replacement and repair, (including upkeep of the computer clusters owned by the group), plus additional funds for miscellaneous supplies and consumables. The breakdown per year is as follows: Year 1 -$4,025, Year 2 -$3,811, Year 3 -$4,042.

Tuition remission for graduate students is a required charge at ASU for graduate students supported by external funding. The remission is charged to projects in proportion to the amount of effort the graduate student will work on the project, and those tuition charges are exempt from Facilities and Administrative Costs. The tuition charge per graduate student is based upon the currently approved university tuition and fee schedules and is calculated as follows: Year 1 - $19,178; Year 2 - $20,712; Year 3 - $22,369.

No participant costs or equipment purchases are budgeted.

8.8 Total direct costs

The total direct costs for each budget year of this award will be $164,264, $171,246, and $179,073 for Years 1, 2, and 3, respectively.

8.9 Indirect costs

Indirect costs are calculated on Modified Total Direct Costs (MTDC) using Facilities and Administrative Costs rates approved by DHHS on August 21, 2018. Items excluded from the indirect cost calculation include are capital equipment, subcontracts over the first $25,000, tuition remission, participant support costs, rental/maintenance of off-campus space, and patient care fees. The off-campus research rate of 26% is applied for all years of the project.

8.10 Total Direct and Indirect costs

The total direct and indirect costs being requested for each 12-month period of this award amount to $197,000, $205,000, and $214,000 for Years 1, 2, and 3, respectively. The total over the requested three-year periods $616,000.

9 Institutional support

Arizona State University provides support to to ASU Meson Physics Group at the Polytechnic campus located in Mesa, AZ. ASU provides utilities, basic telecommunication service, 1 Gbs internet access, approximately 500 square feet of dedicated lab space, and sufficient office space for the principal investigator. These institutional resources are sufficient to support the research efforts related to this proposal for all group members with the proposal funded at the level requested.

10 Summary of meaningful metrics

Metrics that are common to each year:

• TPOL (GlueX)

– Provide polarization values for data taken during previous year

• Search for strangeness (CLAS12)

– The PI will continue to assist in analysis issues that arise regarding strangeness production using CLAS12 and will help coordinate the efforts of the Very Stra nge Group.

– Undergraduate students will obtain experience in analyzing nuclear physics data

by reconstructing various strangeness states.

Metrics specific to year 1:

• Meson resonance analysis (GlueX) – Alan Gardner will obtain PWA results for

• Cascade analysis (GlueX and CLAS12)

– Metrics specific to year 2:

• Meson resonance analysis (GlueX)

– Alan Gardner will have pushed his PWA to include mass states as far above 2 GeV as is reaso nably possible.

– Alan Gardner will provide his prospectus and successfully complete all other re-

quirements to obtain the status of PhD candidate at Arizona State University.

• Cascade analysis (GlueX and CLAS12) Metrics specific to year 3:

• Meson resonance analysis (GlueX)

– We will finish our work on the exploration of meson resonances with K^{+} K^{−} π^{0} decay products

– Alan Gardner will provide the GlueX collaboration with an analysis note regarding his studies and, ideally, lead in the publication of the results.

– Alan Gardner will have successfully defended his dissertation and will have re-

ceived his Ph.D.

• Cascade analysis (GlueX and CLAS12)

Appendices

A Data management plan

The ASU Meson Physics Group, as a member of the CLAS Collaboration and the GlueX Collaboration, adheres to the data management practices of those groups, which are consistent with best practices in the field. As is common in experimental hadronic physics, the data obtained from experiments pass through three stages: raw data, calibrated (or “cooked”) data, and final results. Data in the first two stages are kept in the storage facilities at Jefferson Lab, where those data can be accessed and utilized by institutions, like ASU, that are in the particular collaboration within which the data were taken. As analyses proceed to the results stage, preliminary results are disseminated in talks and presentations. With final collaboration approval, the results are published in refereed journals and placed in collaboration archives for public access.

Expected data: Raw and cooked data are stored physically at JLab. Those JLab storage resources are multiply redundant in order to prevent any data loss. As is the case for other groups, the ASU Meson Physics Group makes copies of portions (or, in some cases, the entirety) of the cooked data for a given running period temporarily using hard drives on computer resources local to the group; that temporary storage also is usually on redundant devices. Thus, at any point in the raw and cooked stages, a given set of data may exist on many other sites than JLab, so preservation is robust. There generally is no attempt by a local group to permanently store cooked data on local resources. In case of local hardware failures or human error leading to the loss of data on local resources, those data are re-copied onto local hardware from the original copies at JLab. Once final results are published, those results are placed on archives accessible to the general public (e.g. Refs. [13,14]).

Data format: The data format for raw and cooked data for the GlueX and CLAS collaborations, called EVIO, was developed by JLab; documentation describing the EVIO format is available in Ref. [15].

Access and data sharing: Raw and cooked data are accessible to members of each collaboration. A person or institution desiring access to the raw and cooked data must apply and be accepted for membership within the CLAS or GlueX collaborations. The final data products of the ASU Meson Physics Group, consisting primarily of cross sections and spin observables, are made publicly available through publications, talks, and presentations, as well as in databases such as the SAID database [14] and CLAS Archive [13].

Policies for re-use and re-distribution: Preliminary results often are presented at open national and international scientific conferences. These preliminary results are internally reviewed within the respective collaboration prior to presentation, with permission obtained to present the preliminary results required by the collaboration prior to the presentation. Such preliminary results are useful in the the initial development of physical interpretations, and comparison with other world data. Reproduction of preliminary results is not permitted without explicit permission of the respective collaborations. Once results have been through the full internal review process within each collaboration, the results are deemed “final”. The data products are submitted to professional journals and made available in public databases such as the SAID database [14] and the CLAS Archive [13].

Archiving of data: As noted above, the raw and cooked data for data runs are archived on tape on-site at JLab. Final results are archived in databases such as the SAID database [14] and the CLAS Archive [13].

B Bibliography and references cited

References

[1] A. Aprahamian et al., Reaching for the Horizon: The 2015 Long Range Plan for Nuclear Science, Nuclear Science Advisory Committee (2015),

URL: {http://science.energy.gov/~/media/np/nsac/pdf/2015LRP/2015_LRPNS_
091815.pdf}

[2] Scheuer, K. and Cole, S. and Dugger, M., Evaluation of machine learning methods on simulated tracks to improve K pi identification in GlueX,

URL:  http://meson.hldsite.com/presentations/Poster2020/
PresentationKevin2020.pdf

[3] G. Collaboration, Mapping the Spectrum of Light Quark Mesons and Gluonic Excitations with Linearly Polarized Photons, Jefferson Lab Experiment Proposal PR-06-102, GlueX Collaboration, Jefferson Lab (2006),

URL: {https://halldweb.jlab.org/DocDB/0012/001226/001/pac30.pdf}

[4] C. Meyers, Potential GlueX Physics During the First Two Years of 12 GeV Running, Report, GlueX Collaboration, Jefferson Lab (2006),

URL: {https://halldweb.jlab.org/DocDB/0006/000621/002/gluex_pac.pdf}

[5] I. Adam et al., The DIRC particle identification system for the BaBar experiment, Nucl. Instrum. Meth. A538 (2005), pp. 281–357

[6] C. Patrignani, Review of Particle Physics, Chin. Phys. C40 (2016) (10), p. 100001

[7] S. Capstick and N. Isgur, Baryons in a relativized quark model with chromodynamics, Phys. Rev. D34 (1986), p. 2809, [AIP Conf. Proc.132,267(1985)]

[8] V. Crede and S. Capstick, Bound States of N’s and Ξ’s, in Workshop on Excited Hyperons in QCD Thermodynamics at Freeze-Out (YSTAR2016) Mini-Proceedings (2017), pp. 40–45,

URL: http://inspirehep.net/record/1511541/files/1510585_40-45.pdf

[9] M. I. Adamovich et al., First observation of the Xi- pi+ decay mode of the Xi(1690)0 hyperon, Eur. Phys. J. C5 (1998), pp. 621–624

[10] J. J. Dudek, R. G. Edwards, P. Guo and C. E. Thomas, Toward the excited isoscalar meson spectrum from lattice QCD, Phys. Rev. D88 (2013) (9), p. 094505 [11] S. U. Chung, E. Klempt and J. G. Korner, SU(3) classification of p−wave ηπ and η′π

systems, Eur. Phys. J. A15 (2002), pp. 539–542

[12] F. Iddir and A. S. Safir, The Decay of the observed J^{P C} = 1^{−+}(1400) and J^{P C} = 1^{−+}(1600) hybrid candidates, Phys. Lett. B507 (2001), pp. 183–192

[13] F. E. Close and H. J. Lipkin, New Experimental Evidence for Four Quark Exotics:The Serpukhov ϕπ Resonance and the Gams ηπ Enhancement, Phys. Lett. B196 (1987), p. 245

[14] CLAS Collaboration, JLAB Experiment CLAS Database,

    URL: http://clasweb.jlab.org/physicsdb/intro.html

[15] Institute for Nuclear Studies Data Analysis Center, SAID Partial-Wave Analysis Facility,

    URL: http://gwdac.phys.gwu.edu/

[16] Jefferson Lab Data Acquisition Group, EVIO Users’ Guide,

    URL: ftp://ftp.jlab.org/pub/coda/evio/2.0/evio_Users_Guide.pdf

C Principal collaborators

C.1 Close collaborators

Hall B/CLAS/CLAS12: B. Briscoe (George Washington University), V. Burkert (JLab), V. Crede (Florida State University), L. Elouadrhiri (JLab), D. Glazier (University of Glasgow), L. Guo (Florida International University), K. Hicks (Ohio University), F. Klein (George Washington University), K. Livingston (University of Glasgow), E. Pasyuk (JLab), J. Price (CSU Dominguez Hills), D. Sober (CUA), I. Strakovsky (George Washington University) S. Strauch (University of South Carolina).

Hall D/GlueX: E. Chudakov (JLab), S. Dobbs (Florida State University), P. Eugenio (Florida State University), R. Jones (University of Connecticut), D. Lawrence (JLab), C.A. Meyer (Carnegie Mellon University), M. Shepherd (Indiana University), E. Smith (JLab), J. Stevens (College of William and Mary), S. Taylor (JLab).

C.2 Advisors and postdoctoral sponsors of the investigators

MRD: B.G. Ritchie (ASU); BGR: F.T. Avignone, III (University of South Carolina), B.M. Preedom (University of South Carolina), P.G. Roos (University of Maryland).

C.3 Former post-doctoral participants in academic and lab positions

E.A. Pasyuk: Senior Staff Scientist, Jefferson Lab; M.R. Dugger: Associate Professor, ASU.

C.4 Former Ph.D. students in academic and lab positions

M. Dugger: Associate Professor, ASU; D. Lawrence: Staff Scientist, JLab; R. Tucker: Lecturer, ASU.

D Biographical sketches

D.1 Michael R. Dugger

Professional preparation

\underline{Institution} \underline{Major/Conc.} \underline{Degree, Year}
Undergraduate: Northern Arizona State University Physics/Mathematics B.S., 1993
Graduate:  Arizona State University  Physics  Ph.D., 2001
Postdoctoral: Arizona State University  Physics  2002-2006

Appointments

Associate Professor (tenure-track) Arizona State University 2017 to present
Associate Research Professor  Arizona State University 2013 to 2017
Assistant Research Professor  Arizona State University 2006-2013
Postdoctoral Research Associate Arizona State University 2002-2006

Invited talks over the last three years

“Overview of Spectroscopy Results in Meson Photoproduction with Polarization Observables,” HADRON, Newport News, VA., Septrmber 14, 2015.

Synergistic activities

GlueX Collaboration Board member, 2013-2015; Division of Nuclear Physics, Conference Experience for Undergraduates, Review Committee member, 2018; International Advisory Committee, MENU 2019 Conference, 2018-Present; CLAS Collaboration, Analysis Coordinator for Ξ and Ω^{−} photoproduction, 2017-Present.

Publications over the last three years

1. M. C. Kunkel et al., Exclusive photoproduction of π^{0} up to large values of Mandelstam variables s, t and u with CLAS, Phys. Rev. C98 (2018) (1), p. 015207

2. P. Roy et al., Measurement of the beam asymmetry Σ and the target asymmetry T in the photoproduction of ω mesons off the proton using CLAS at Jefferson Laboratory, Phys. Rev. C97 (2018) (5), p. 055202

3. Z. Akbar et al., Measurement of the helicity asymmetry E in ω → π^{+}π^{−}π^{0} photoproduction, Phys. Rev. C96 (2017) (6), p. 065209

4. A. V. Anisovich, V. Burkert, P. M. Collins, M. Dugger, E. Klempt, V. A. Nikonov, B. G. Ritchie, A. V. Sarantsev and U. Thoma, N^{∗} → Nη′ decays from photoproduction of η′-mesons off protons, Phys. Lett. B772 (2017), pp. 247–252

5. P. Collins et al., Photon beam asymmetry Σ in the reaction γ⃗p → pω for E_{γ} = 1.152 to 1.876 GeV, Phys. Lett. B773 (2017), pp. 112–120

6. P. T. Mattione et al., Differential cross section measurements for γn → π^{−}p above the first nucleon resonance region, Phys. Rev. C96 (2017) (3), p. 035204 7. A. V. Anisovich et al., Differential cross sections and polarization observables from CLAS

K* photoproduction and the search for new N* states, Phys. Lett. B771 (2017), pp. 142–150

8. M. Dugger et al., Design and construction of a high-energy photon polarimeter, Nucl. Instrum. Meth. A867 (2017), pp. 115–127

9. P. Collins et al., Photon beam asymmetry Σ for η and η′ photoproduction from the proton, Phys. Lett. B771 (2017), pp. 213–221

10. H. Al Ghoul et al., Measurement of the beam asymmetry Σ for π^{0} and η photoproduction on the proton at E_{γ} = 9 GeV, Phys. Rev. C95 (2017) (4), p. 042201

11. C. A. Paterson et al., Photoproduction of Λ and Σ^{0} hyperons using linearly polarized photons, Phys. Rev. C93 (2016) (6), p. 065201

12. H. Al Ghoul et al., First results from the GlueX experiment, AIP Conf. Proc. 1735 (2016), p. 020001